### v2

# Commute Graph Neural Networks

Abstract
Graph Neural Networks (GNNs) have shown remarkable success in learning from graph-structured data. However, their application to directed graphs (digraphs) presents unique challenges, primarily due to the inherent asymmetry in node relationships. Traditional GNNs are adept at capturing unidirectional relations but fall short in encoding the mutual path dependencies between nodes, such as asymmetrical shortest paths typically found in digraphs. Recognizing this gap, we introduce Commute Graph Neural Networks (CGNN), an approach that seamlessly integrates node-wise commute time into the message passing scheme. The cornerstone of CGNN is an efficient method for computing commute time using a newly formulated digraph Laplacian. Commute time is then integrated into the neighborhood aggregation process, with neighbor contributions weighted according to their respective commute time to the central node in each layer. It enables CGNN to directly capture the mutual, asymmetric relationships in digraphs. Extensive experiments confirm the superior performance of CGNN.
1 Introduction
Directed graphs (digraphs) are widely employed to model relational structures in diverse domains, such as social networks [1] and recommendation systems [2]. Recently, the advances of graph neural networks (GNNs) have inspired various attempts to adopt GNNs for analyzing digraphs [3, 4, 5, 6, 7, 8]. The essence of GNN-based digraph analysis lies in utilizing GNNs to learn expressive node representations that encode edge direction information.
To achieve this, modern digraph neural networks are designed to integrate edge direction information into the message passing process by distinguishing between incoming and outgoing edges. This distinction enables the central node to learn directionally discriminative information from its neighbors. As illustrated in the digraph of Fig. 1, given a central node $v_{i}$, a 1-layer digraph neural network can aggregate messages from $v_{i}$’s incoming neighbor $v_{m}$ and outgoing neighbor $v_{j}$, and simultaneously capture edge directions by applying direction-specific aggregation functions [7], or by predefining edge-specific weights [6, 4].
Figure 1: A digraph and its undirected counterpart. Blue arrows indicate unidirectional paths, together with longer paths in the gray area, forming commute closed loops between the central node $v_{i}$ and its outgoing neighbors $v_{j}$ and $v_{k}$. In the undirected graph, shortest path distances (SPD) between nodes are symmetric. However, in the digraph, the fact that unidirectional SPDs are equal does not imply that mutual SPDs will also be equal. For instance, while the SPDs from $v_{i}$ to $v_{j}$ and $v_{k}$ are identical, the reverse SPD from $v_{j}$ and $v_{k}$ back to $v_{i}$ do not necessarily match these distances.
Despite the advancements, current digraph neural networks primarily capture unidirectional111‘unidirectional’ refers to relationships in digraphs where edges have a specific direction from one node to another. relationships between nodes, neglecting the complexity arising from path asymmetry. For instance, a $k$-layer GNN aggregates the neighbors within the shortest path $k$ for the central node. If the graph is undirected, the shortest path between any two nodes is symmetric, as shown in the undirected graph of Fig. 1. This symmetry simplifies the representation of node relationships, implying that if the SPDs from one node to two other nodes are identical, then the SPDs from these two nodes back to the source node must also be the same. Conversely, such symmetry is absent in digraphs. Considering the digraph in Fig. 1, the shortest paths between $v_{i}$ and $v_{j}$ are asymmetric. Therefore, although $v_{j}$ and $v_{k}$ are both immediate outgoing neighbors of $v_{i}$, the strength of their relationships with the central node differs significantly. Existing methods [7, 4, 6], by focusing solely on unidirectional shortest paths (blue and red arrows), fail to capture the asymmetry phenomenon, which conveys valuable information of node relationships. Take social networks as an example: an ordinary user can directly follow a celebrity, yielding a short path to the celebrity, yet the reverse path from the celebrity to the follower might be much longer. Considering only the short path from the follower to celebrity could falsely suggest a level of closeness that does not exist. In contrast, accounting for the mutual paths between users yields a more precise and robust measure of their relationship, with stronger mutual interactions implying stronger connections.
To capture the mutual path interactions in GNNs, we adapt the concept of commute time, the expected number of steps to traverse from a source node to a target and back, from the Markov chain theory to the domain of graph learning. To this end, we first generalize the graph Laplacian to the digraph by defining the divergence of the gradient on the digraph. Utilizing this digraph-specific Laplacian, we develop an efficient method to compute commute time, ensuring sparsity and computational feasibility. Then we incorporate the commute-time-based proximity measure into the message passing process by assigning aggregation weights to neighbors. The intuition behind is that the immediate and unidirectional neighboring relationships do not necessarily imply strong similarity, but the mutual proximity is a more reliable indicator of relationship closeness. Our experimental results demonstrate the effectiveness of CGNN.
2 Related Work
2.1 Digraph Laplacian
While the Laplacian for undirected graphs has been extensively studied, the area of Laplace operator digraphs remains underexplored. [9] pioneers this area by defining a normalized Laplace operator specifically for strongly connected directed graphs with nonnegative weights. This operator is expressed as $\mathbf{I}-\frac{\pi^{1/2}\mathbf{P}\pi^{-1/2}+\pi^{-1/2}\mathbf{P}^{*}\pi^{1/2}}{2}$. Key to this formulation is the use of the transition probability operator $\mathbf{P}$ and the Perron vector $\pi$, with the operator being self-adjoint. Building on the undirected graph Laplacian, [10] adapt this concept to accommodate the directed structure, focusing particularly on the in-degree matrix. They define the directed graph Laplacian as $\mathbf{D}_{\text{in}}-\mathbf{A}$, where $\mathbf{D}_{\text{in}}=\operatorname{diag}\left(\left\{d_{i}^{\text{in}}\right\}_{i=1}^{N}\right)$ represents the in-degree matrix. [11] uses stationary probabilities of the Markov chain governing random walks on digraphs to define the Laplacian as $\pi^{\frac{1}{2}}(\mathbf{I}-\mathbf{P})\pi^{-\frac{1}{2}}$, which underscores the importance of random walks and their stationary distributions in understanding digraph dynamics. Hermitian Laplacian [12] consider the edge directionality and node connectivity separately, and encode the edge direction into the argument in the complex plane. Diverging from existing Laplacians, our proposed DiLap $\mathbf{\Pi}(\mathbf{D}^{-1}-\mathbf{P})$ is grounded in graph signal processing principles, conceptualized as the divergence of a signal’s gradient on the digraph. It encompasses the degree matrix $\mathbf{D}$ to preserve local connectivity, the transition matrix $\mathbf{P}$ to maintain the graph’s directed structure, and the diagonalized Perron vector $\mathbf{\Pi}$, capturing critical global graph attributes such as node structural importance, global connectivity, and expected reachability [13].
2.2 Digraph Neural Networks
To effectively capture the directed structure with GNNs, spectral-based methods [6, 3, 4] have been proposed to preserve the underlying spectral properties of the digraph by performing spectral analysis based on the digraph Laplacian proposed by [9]. [14] introduce holomorphic filters as spectral filters for digraphs, and investigate their optimal filter-bank. MagNet [6] and its extensions [15, 16] utilizes magnetic Laplacian to derive a complex-valued Hermitian matrix to encode the asymmetric nature of digraphs. Spatial GNNs also offer a natural approach to capturing directed structures. For instance, GraphSAGE [17] allows for controlling the direction of information flow by considering in-neighbors or out-neighbors separately. DirGNN [7] further extends this framework by segregating neighbor aggregation according to edge directions, offering a more refined method to handle the directed nature of graphs. Transformer-based methods capture directed structure by specific positional encoding modules, such as directional random walk encoding [8] and partial order encoding [18]. DUPLEX [19] utilizes Hermitian adjacency matrix decomposition for neighbor aggregation and incorporates a dual GAT encoder for modeling directional neighbors.
3 Preliminary
Notations
Consider $G=(V,E,\mathbf{X})$ as an unweighted digraph comprising $N$ nodes, where $V=\{v_{i}\}^{N}_{i=1}$ is the node set, $E\subseteq(V\times V)$ is the edge set, $\mathbf{X}\in\mathbb{R}^{N\times d}$ is the node feature matrix. $Y=\{y_{1},\cdots,y_{N}\}$ is the set of labels for $V$. Let $\mathbf{A}\in\mathbb{R}^{N\times N}$ be the adjacency matrix and $\mathbf{D}=\mathrm{diag}(d_{1},\cdots,d_{N})\in\mathbb{R}^{N\times N}$ be the degree matrix of $\mathbf{A}$, where $d_{i}=\sum_{v_{j}\in V}\mathbf{A}_{ij}$ is the out-degree of $v_{i}$. Let $\tilde{\mathbf{A}}=\mathbf{A}+\mathbf{I}$ and $\tilde{\mathbf{D}}=\mathbf{D}+\mathbf{I}$ denote the adjacency and degree matrix with self-loops, respectively. The transition probability matrix of the Markov chain associated with random walks on $G$ is defined as $\mathbf{P}=\mathbf{D}^{-1}\mathbf{A}$, where $\mathbf{P}_{ij}=\mathbf{A}_{ij}/\mathrm{deg}(v_{i})$ is the probability of a 1-step random walk starting from $v_{i}$ to $v_{j}$. Graph Laplacian formalized as $\mathbf{L}=\mathbf{D}-\mathbf{A}$ is defined on the undirected graph whose adjacency matrix is symmetric. The symmetrically normalized Laplacian with self-loops [20] is defined as $\hat{\mathbf{L}}=\tilde{\mathbf{D}}^{-\frac{1}{2}}\tilde{\mathbf{L}}\tilde{\mathbf{D}}^{-\frac{1}{2}}$, where $\tilde{\mathbf{L}}=\tilde{\mathbf{D}}-\tilde{\mathbf{A}}$.
Digraph Neural Networks
DirGNN [7] is a general framework that generalizes the message passing paradigm to digraphs by adapting to the directionality of edges. It involves separate aggregation processes for incoming and outgoing neighbors of each node as follows:
$m_{i,\text{in}}^{(\ell)}$ $=\mathrm{Agg}_{\text{in}}^{(\ell)}\left(\left\{h_{j}^{(\ell-1)}:v_{j}\in\mathcal{N}_{i}^{\text{in}}\right\}\right)$ (1)
$m_{i,\text{out}}^{(\ell)}$ $=\mathrm{Agg}_{\text{out}}^{(\ell)}\left(\left\{h_{j}^{(\ell-1)}:v_{j}\in\mathcal{N}_{i}^{\text{out}}\right\}\right)$
$h_{i}^{(l)}$ $=\mathrm{Comb}^{(\ell)}\left(h_{i}^{(\ell-1)},m_{i,\text{in}}^{(\ell)},m_{i,\text{out}}^{(\ell)}\right),$
where $\mathcal{N}_{i}^{\text{in}}$ and $\mathcal{N}_{i}^{\text{out}}$ are respectively incoming and outgoing neighbors of $v_{i}$. $\mathrm{Agg}_{\text{in}}^{(\ell)}(\cdot)$ and $\mathrm{Agg}_{\text{out}}^{(\ell)}(\cdot)$ are specialized aggregation functions of $\mathcal{N}_{i}^{\text{in}}$ and $\mathcal{N}_{i}^{\text{out}}$ at layer $\ell$, used to encode the directional characteristics of the edges connected to $v_{i}$.
4 Random Walk Distance and GNNs
Based on the established notations. We then show that message passing based GNNs naturally capture the concept of hitting time during information propagation across the graph, due to the unidirectional nature of the neighborhood aggregation. Subsequently, we argue for the significance of commute time, highlighting it as a more compact measure of mutual node-wise interactions in random walks.
4.1 Can GNNs Capture Random Walk Distance?
In the context of random walks on a digraph, hitting time and commute time, collectively referred to as random walk distances, serve as key metrics for assessing node connectivity and interaction strength. Hitting time $h(v_{i},v_{j})$ is the expected number of steps a random walk takes to reach a specific target node $v_{j}$ for the first time, starting from a given source node $v_{i}$. Commute time $c(v_{i},v_{j})$ is the expected number of steps required for a random walk to start at $v_{i}$, reach $v_{j}$, and come back. A high hitting (commute) time indicates difficulty in achieving unidirectional (mutual) visits to each other in a random walk. As illustrated in the digraph of Fig. 1, commute time $c(v_{i},v_{j})>c(v_{i},v_{k})$, while the hitting time $h(v_{m},v_{i})=h(v_{i},v_{j})=h(v_{i},v_{k})$.
Motivation
Given these definitions, two questions arise: How crucial is it to retain these measures in graph learning? Also, are message-passing GNNs capable of preserving these characteristics? Firstly, both hitting time and commute time are critical in understanding the structural dynamics of graphs. Hitting time, analogous to the shortest path, measures the cost of reaching one node from another, reflecting the directed influence or connectivity. Commute time, encompassing the round-trip journey, offers insights into the mutual relationships between nodes, which is especially evident in social networks, as illustrated by celebrity-follower relationships. Secondly, message-passing GNNs are somewhat effective in capturing hitting time, as they propagate information across the graph in a manner similar to a random walk, where quickly reached nodes are preferentially aggregated, and the influence of nodes exponentially diminishes with increasing distance [21]. However, GNNs face challenges in preserving commute time due to their requirement for comprehending mutual path relations, which are inherently asymmetric and often involve longer-range interactions especially in digraphs, which are not naturally captured in the basic message-passing framework.
Taking the digraph in Fig. 1 as an example, a 1-layer DirGNN defined in Eq. 1 can encode $v_{m}$, $v_{j}$ and $v_{k}$ into the representation of $v_{i}$, while also capturing the directionality of edges from these neighbors by using distinct aggregation functions for incoming and outgoing neighbors. It shows that DirGNN can capture the hitting time, as neighbors with lower hitting times, $h(v_{i},v_{k})$, $h(v_{i},v_{k})$ and $h(v_{m},v_{i})$, are aggregated preferentially. However, DirGNN inherently focuses on unidirectional interactions and overlooks mutual path dependencies. Notably, a 1-layer DirGNN is insufficient for capturing the asymmetric interactions indicated by the paths from $v_{j}$ and $v_{k}$ returning to the central node $v_{i}$ (gray areas). One potential approach to address this limitation is to stack additional message passing layers to encompass the entire commute path between nodes, thereby capturing mutual path interactions. Nevertheless, this strategy is non-trivial because the commute paths vary considerably across different node pairs, complicating the determination of an appropriate number of layers. Additionally, stacking multiple layers to cover these paths can introduce irrelevant non-local information and lead to oversmoothing.
Goal
We expect to directly encode node-wise commute time into the node representations to accurately reflect the true interaction strength between adjacent nodes during neighbor aggregation, accounting for both forward and backward paths. For instance, even though $h(v_{i},v_{j})=h(v_{i},v_{k})$, a shorter commute time $c(v_{i},v_{k})<c(v_{i},v_{j})$ suggests a stronger interaction from $v_{k}$ to $v_{i}$ compared to $v_{j}$ to $v_{i}$. Consequently, the contribution of neighbor $v_{k}$ to the representation of $v_{i}$ should be greater than that of $v_{j}$.
4.2 Commute Time Computation
Based on the standard Markov chain theory, a useful tool to study random walk distances is the fundamental matrix [22]. We first establish the following assumptions required to support the theorem.
Assumption 4.1.
The digraph $G$ is irreducible and aperiodic.
These two properties pertain to the Markov chain’s stationary probability distribution $\pi$ (i.e., Perron vector) corresponding to the given graph. Irreducibility ensures that it is possible to reach any node (state) from any other node, preventing $\pi$ from converging to 0. Aperiodicity ensures that the Markov chain does not get trapped in cycles of a fixed length, thus guaranteeing the existence of a unique $\pi$. Existence and uniqueness of $\pi$ facilitate deterministic analysis and computation. For a more intuitive understanding of the assumptions, we give the sufficient conditions of digraph under the irreducibility and aperiodicity assumptions.
Proposition 1.
A strongly connected digraph, in which a directed path exists between every pair of vertices, is irreducible. A digraph with self-loops in each node is aperiodic.
Given the above assumption, the fundamental matrix $\mathbf{Z}$ is defined as the sum of an infinite matrix series:
$\mathbf{Z}=\sum_{t=0}^{\infty}\left(\mathbf{P}^{t}-\mathbf{J}\mathbf{\Pi}\right)=\sum_{t=0}^{\infty}\left(\mathbf{P}^{t}-e\pi^{\top}\right),$ (2)
where $e$ is the all-one column vector, then we have $\mathbf{J}=e\cdot e^{\top}$ is the all-one matrix, and $\mathbf{\Pi}=\mathrm{diag}(\pi)$ is the diagonal matrix of $\pi$.
Theorem 4.2.
[11] Given 4.1, the fundamental matrix $\mathbf{Z}$ defined in Eq. 2 converges to:
$\mathbf{Z}=(\mathbf{I}-\mathbf{P}+\mathbf{J}\mathbf{\Pi})^{-1}-\mathbf{J}\mathbf{\Pi},$ (3)
where $\mathbf{I}$ is an identity matrix.
The hitting time and commute time on $G$ can then be expressed as $\mathbf{Z}$ [22] as follows:
$h(v_{i},v_{j})=\frac{\mathbf{Z}_{jj}-\mathbf{Z}_{ij}}{\pi_{j}},\quad c(v_{i},v_{j})=h(v_{i},v_{j})+h(v_{j},v_{i}).$ (4)
However, directly calculating the complete fundamental matrix $\mathbf{Z}$ and the commute times for all node pairs is computationally expensive and yields a dense matrix. Moreover, integrating the random walk distances computation, defined in Eq. 3 and Eq. 4, into the message passing framework is non-trivial, which concerns the scalability of the model.
5 Commute Graph Neural Networks
In this section, we present Commute Graph Neural Networks (CGNN) to encode the commute time information into message passing. We first establish the relationship between random walk distances and the digraph Laplacian.
5.1 Digraph Laplacian (DiLap)
Contrary to the traditional graph Laplacian, typically defined as a symmetric positive semi-definite matrix derived from the symmetric adjacency matrix, our proposed DiLap is built upon the transition matrix to preserve the directed structure. Specifically, the classical graph Laplacian $\mathbf{L}=\mathbf{D}-\mathbf{A}$ is interpreted as the divergence of the gradient of a signal on an undirected graph [23, 24]: given a graph signal $s\in\mathbb{R}^{N}$, $(\mathbf{L}s)_{i}=\sum_{j\in\mathcal{N}_{i}}\mathbf{A}_{ij}(s_{i}-s_{j})$. Intuitively, graph Laplacian corresponds to the difference operator on the signal $s$, and acts as a node-wise measure of local smoothness. In line with this conceptual foundation, we generalize the graph Laplacian to digraphs by defining DiLap $\mathbf{T}$:
$(\mathbf{T}s)_{i}=\sum_{v_{j}\in\mathcal{N}^{\text{out}}_{i}}\mathbf{P}_{ij}(s_{i}-s_{j})=((\mathbf{D}^{-1}-\mathbf{P})s)_{i},\quad\text{wehave}\quad\mathbf{T}=\mathbf{D}^{-1}-\mathbf{P},$ (5)
where $\mathcal{N}^{\text{out}}_{i}$ is the set of $v_{i}$’s out-neighbors, and $\mathbf{P}_{ij}=\frac{1}{d^{\text{out}}_{i}}$ with $d^{\text{out}}_{i}$ being the out-degree of $v_{i}$. The detailed derivation of $\mathbf{T}$ is included in Section 8.1. Eq. 5 illustrates that DiLap $\mathbf{T}$ acts as the divergence of the gradient of a given signal on a digraph, and can be used to measure the local smoothness in a digraph.
Considering the Laplacian operator’s function in measuring the smoothness of a signal across the entire graph, it is crucial to assign greater weights to nodes of higher structural importance. It is because the smoothness at nodes central to the graph structure should have a more pronounced impact on the global smoothness measurement. Thus, we further define the Weighted DiLap $\mathcal{T}$:
$(\mathcal{T}s)_{i}=\sum_{v_{j}\in\mathcal{N}^{\text{out}}_{i}}\pi_{i}\mathbf{P}_{ij}(s_{i}-s_{j})=(\mathbf{\Pi}(\mathbf{D}^{-1}-\mathbf{P})s)_{i},\quad\text{we have}\quad\mathcal{T}=\mathbf{\Pi}(\mathbf{D}^{-1}-\mathbf{P}).$ (6)
Here we utilize the $i$-th element of the Perron vector $\pi$ to quantify the structural importance of $v_{i}$, reflecting its eigenvector centrality. This is based on the principle that a node’s reachability is directly proportional to its corresponding value in the Perron vector [25]. Therefore, $\pi$ effectively indicates the centrality and influence over the long term in the graph. Perron-Frobenius Theorem [26] establishes that $\pi$ satisfies $\sum_{i}\pi(i)=1$, is strictly positive, and converges to the left eigenvector of the dominant eigenvalue of $\mathbf{P}$.
5.2 Similarity-based Graph Rewiring
Both the fundamental matrix defined in Eq. 3 and Weighted DiLap requires 4.1 to ensure the existence and uniqueness of the Perron vector $\pi$, conditions that are not universally met in general graphs. To fulfill the irreducibility and aperiodicity assumptions,  [3] introduce a teleporting probability uniformly distributed across all nodes. This method, inspired by PageRank [27], amends the transition matrix to $\mathbf{P}_{pr}=\gamma\mathbf{P}+(1-\gamma)\frac{ee^{\top}}{N}$, where $\gamma\in(0,1)$. $\mathbf{P}_{pr}$ allows for the possibility that a random walker may choose a non-neighbor node for the next step with a probability of $\frac{1-\gamma}{N}$. This adjustment ensures that $\mathbf{P}_{pr}$ is irreducible and aperiodic, so it has a unique $\pi$. However, this approach leads to a complete graph represented by a dense matrix $\mathbf{P}_{pr}$, posing significant challenges for subsequent computational processes.
Rather than employing $\mathbf{P}_{pr}$ as the transition matrix, we introduce a graph rewiring method based on feature similarity to make a given graph irreducible, while maintaining the sparsity. As outlined in Proposition 1, to transform the digraph into a strongly connected structure, it is essential that each node possesses a directed path to every other node. To this end, we initially construct a simple and irreducible graph $G^{\prime}$ with all $N$ nodes, then add all edges from $G^{\prime}$ the original digraph $G$, thereby ensuring $G$’s irreducibility. The construction of $G^{\prime}$ begins with the calculation of the mean of node features as the anchor vector $\bm{a}$. Then we determine the similarity between each node and the anchor, sort the similarity values, and return the sorted node indices, denoted as $S\in\mathbb{R}^{N}$:
$\bm{a}=\frac{\sum_{i}\mathbf{X}_{i}}{N},\quad s_{i}=\cos(\bm{a},\mathbf{X}_{i}),\quad S=\arg\mathrm{sort}(\{s_{i}\}_{i=1}^{N})$ (7)
where $\cos(\bm{a},\mathbf{X}_{i})$ is the cosine similarity between node features of $v_{i}$ and $\bm{a}$, and $\mathrm{argsort}(\cdot)$ yields the indices of nodes that sort similarity values $\{s_{i}\}_{i=1}^{N}$. We then connect the nodes one by one with undirected (bidirectional) edges following the order in $S$ to construct $G^{\prime}$, as shown in Fig. 2. Given that $G^{\prime}$ is strongly connected, adding all its edges into $G$ results in a strongly connected digraph $\widetilde{G}$, which is irreducible. To achieve aperiodicity, self-loops are further added to $\widetilde{G}$.
Figure 2: The sorted node indices in $S$ are connected one by one with undirected edges to construct $G^{\prime}$, then adding all edges from $G^{\prime}$ to $G$ to generate $\widetilde{G}$.
This rewiring approach satisfies 4.1 and maintains graph sparsity. Additionally, adding edges between nodes with similar features only minimally alters the overall semantics of the original graph. Based on $\widetilde{G}$ and its corresponding $\widetilde{\mathbf{P}}$ and $\widetilde{\mathbf{\Pi}}$, we have the deterministic Weighted DiLap $\widetilde{\mathcal{T}}$.
5.3 From DiLap to Commute Time
Given the Weighted DiLap $\widetilde{\mathcal{T}}$, we can unify the commute time information into the message passing by building the connection between $\widetilde{\mathcal{T}}$ and the fundamental matrix $\mathbf{Z}$:
Lemma 5.1.
Given a rewired graph $\widetilde{G}$, the Weighted DiLap is defined as $\widetilde{\mathcal{T}}=\widetilde{\mathbf{\Pi}}(\widetilde{\mathbf{D}}^{-1}-\widetilde{\mathbf{P}})$. Then the fundamental matrix $\mathbf{Z}$ of $\widetilde{G}$ can be solved by:
$\mathbf{Z}=\widetilde{\mathbf{\Pi}}^{-\frac{1}{2}}\mathcal{R}^{\dagger}\widetilde{\mathbf{\Pi}}^{\frac{1}{2}},$ (8)
where $\mathcal{R}=\widetilde{\mathbf{\Pi}}^{-\frac{1}{2}}\widetilde{\mathcal{T}}\widetilde{\mathbf{\Pi}}^{-\frac{1}{2}}-\widetilde{\mathbf{D}}^{-1}+\mathbf{I}$ and the superscript $\dagger$ means Moore–Penrose pseudoinverse of the matrix.
The proof is given in Section 8.2. Leveraging Lemma 5.1 and using Eq. 4, we can further compute the hitting times and commute times in terms of $\widetilde{\mathcal{T}}$ with the following theorem.
Theorem 5.2.
Given $\widetilde{G}$, the hitting time and commute time from $v_{i}$ to $v_{j}$ on $\widetilde{G}$ can be computed as follows:
$h(v_{i},v_{j})$ $=\frac{\mathcal{R}^{\dagger}_{jj}}{\pi_{j}}-\frac{\mathcal{R}^{\dagger}_{ij}}{\sqrt{\pi_{i}\pi_{j}}},$ (9)
$c(v_{i},v_{j})$ $=h(v_{i},v_{j})+h(v_{j},v_{i})=\frac{\mathcal{R}^{\dagger}_{jj}}{\pi_{j}}+\frac{\mathcal{R}^{\dagger}_{ii}}{\pi_{i}}-\frac{\mathcal{R}^{\dagger}_{ij}}{\sqrt{\pi_{i}\pi_{j}}}-\frac{\mathcal{R}^{\dagger}_{ji}}{\sqrt{\pi_{i}\pi_{j}}}.$
Then we can derive the matrix forms of the hitting time $\mathcal{H}$ and commute time $\mathcal{C}$ as per Eq. 9:
$\mathcal{H}=(e\otimes\pi^{-1})(\mathcal{R}^{\dagger}\odot\mathbf{I})-\mathcal{R}^{\dagger}\odot(\pi^{-\frac{1}{2}}\otimes\pi^{-\frac{1}{2}}),\quad\mathcal{C}=\mathcal{H}+\mathcal{H}^{\top}$ (10)
where $\odot$ denotes Hadamard product, and $\otimes$ is outer product. $\mathcal{H}_{ij}$ and $\mathcal{C}_{ij}$ correspond to the hitting and commute time from $v_{i}$ to $v_{j}$ respectively. The computation of commute times via DiLap, in contrast to the method delineated in Theorem 4.2, is primarily motivated by efficiency concerns. Specifically, Eq. 3 necessitates the inversion of a dense matrix with complexity $\mathcal{O}(N^{3})$, whereas our DiLap-based method hinges on computing the pseudoinverse of a sparse matrix $\mathcal{R}$. The pseudoinverse of $\mathcal{R}$ can be efficiently determined using SVD. Given the sparse nature of $\mathcal{R}$, we can employ well-established techniques such as the randomized truncated SVD algorithm [28, 29], which takes advantage of sparsity, to reduce the time complexity to $\mathcal{O}(q|E|)$, where $|E|$ denotes the number of edges reflecting the sparsity (See Section 8.3). Next, we present Commute Graph Neural Networks (CGNN) based on $\mathcal{C}$.
5.4 CGNN
$\mathcal{C}\in\mathbb{R}^{N\times N}$ quantifies the strength of mutual relations between node pairs in the random walk context. Notably, smaller values in $\mathcal{C}$ correspond to stronger mutual reachability, indicating stronger relations between node pairs. Thus, $\mathcal{C}$ is a positive symmetric matrix, and the commute-time-based node proximity matrix can be expressed as $\widetilde{\mathcal{C}}=\exp(-\mathcal{C})$. Since the directed adjacency matrix $\mathbf{A}$ represents the outgoing edges of each node, $\mathbf{A}^{\top}$ therefore accounts for all incoming edges. Then we have $\widetilde{\mathcal{C}}^{\text{out}}=\mathbf{A}\odot\widetilde{\mathcal{C}}$ and $\widetilde{\mathcal{C}}^{\text{in}}=\mathbf{A}^{\top}\odot\widetilde{\mathcal{C}}$ represent the proximity between adjacent nodes under outgoing and incoming edges, respectively. We further perform row-wise max-normalization on $\widetilde{\mathcal{C}}^{\text{out}}$ and $\widetilde{\mathcal{C}}^{\text{in}}$ to rescale the maximum value in each row to 1. Given the original graph $G$ as input, we define the $\ell$-th layer of CGNN as:
$m_{i,\text{in}}^{(\ell)}$ $=\mathrm{Agg}_{\text{in}}^{(\ell)}\left(\left\{\widetilde{\mathcal{C}}^{\text{in}}_{ij}\cdot h_{j}^{(\ell-1)}:v_{j}\in\mathcal{N}_{i}^{\text{in}}\right\}\right)$ (11)
$m_{i,\text{out}}^{(\ell)}$ $=\mathrm{Agg}_{\text{out}}^{(\ell)}\left(\left\{\widetilde{\mathcal{C}}^{\text{out}}_{ij}\cdot h_{j}^{(\ell-1)}:v_{j}\in\mathcal{N}_{i}^{\text{out}}\right\}\right)$
$h_{i}^{(l)}$ $=\mathrm{Comb}^{(\ell)}\left(h_{i}^{(\ell-1)},m_{i,\text{in}}^{(\ell)},m_{i,\text{out}}^{(\ell)}\right),$
where $\mathrm{Agg}_{\text{in}}^{(\ell)}(\cdot)$ and $\mathrm{Agg}_{\text{out}}^{(\ell)}(\cdot)$ are mean aggregation functions with different feature transformation weights, and $\mathrm{Comb}^{(\ell)}(\cdot)$ is a mean operator. Within each layer, the influence of $v_{j}$ on the central node $v_{i}$ is modulated by the commute-time-based proximity $\widetilde{\mathcal{C}}$ based on the edge directionality. We present the pseudocode of CGNN in Algorithm 1.
Algorithm 1 CGNN
Input: Digraph $G=(V,E,\mathbf{X})$; Depth $L$; Hidden size $d^{\prime}$; Number of classes $K$
Output: Logits $\hat{Y}\in\mathbb{R}^{N\times K}$
1:  Compute the anchor $\bm{a}$ and node-anchor similarities to construct $G^{\prime}$ with Eq. 7.
2:  Add all edges from $G^{\prime}$ to $G$ to generate $\widetilde{G}$.
3:  Compute the Weight DiLap $\widetilde{\mathcal{T}}$ for $\widetilde{G}$ with Eq. 6.
4:  Compute $\mathcal{R}$ and its Moore-Penrose pseudoinverse with Eq. 8 and Eq. 19.
5:  Compute the commute time matrix $\mathcal{C}$ with Eq. 10.
6:  Compute the normalized proximity matrix $\widetilde{\mathcal{C}}$ with $\widetilde{\mathcal{C}}^{\text{out}}=\mathbf{A}\odot\widetilde{\mathcal{C}}$ and $\widetilde{\mathcal{C}}^{\text{in}}=\mathbf{A}^{\top}\odot\widetilde{\mathcal{C}}$.
7:  for $\ell\in\{1,\cdots,L\}$ do
8:     Layer-wise message passing with Eq. 11.
9:  end for
10:  $\mathbf{H}=\mathrm{MLP}(\mathbf{H}^{(L)})$.
11:  $\hat{Y}=\mathrm{Softmax}(\mathbf{H})$.
Complexity Analysis
The randomized truncated SVD to compute $\mathcal{R}^{\dagger}$ is $\mathcal{O}(q|E|)$ where $q$ is the required rank, and the message passing iteration has the same time complexity as DirGNN with $\mathcal{O}(L|E|d^{2})$. Therefore, the overall time complexity of CGNN is $\mathcal{O}((Ld^{2}+q)|E|)$. In practice, $q$ is typically set to 5, rendering the time complexity effectively linear with respect to the number of edges $|E|$. In GNN domain, models with a complexity less than $\mathcal{O}(N^{2})$ are generally considered feasible by researchers [30]. Given that real-world networks are often extremely sparse, i.e., $|E|\ll N^{2}$, CGNN demonstrates its feasibility as a model within the GNN family.
6 Experiments
6.1 Experimental Settings
We conduct node classification experiments on eight digraph datasets. We provide a performance comparison with 12 baselines including 1) General GNNs: GCN [31], GAT [32], and GraphSAGE [17]; 2) Non-local GNNs: APPNP [33], MixHop [34], GPRGNN [35], and GCNII [36]; 3) Digraph NNs: DGCN [4], DiGCN [3], MagNet [6], DiGCL [5], DUPLEX [19], and DirGNN [7]. For all baselines, we apply both the symmetrized and asymmetric adjacency matrix for experiments. The results reported are the better of the two results.
We evaluate the performance by node classification accuracy with standard deviation in the semi-supervised setting. For Squirrel and Chameleon, we use 10 public splits (48%/32%/20% for training/validation/testing) provided by [37]. For the remaining datasets, we adopt the same splits as [3, 5], which chooses 20 nodes per class for the training set, 500 for the validation set, and allocates the rest to the test set. We conduct our experiments on 2 Intel Xeon Gold 5215 CPUs and 1 NVIDIA GeForce RTX 3090 GPU.
6.2 Data Statistics
The datasets used in Section 6 are Squirrel, Chameleon [38], Citeseer [39], CoraML [40], AM-Photo [41], Snap-Patents, Roman-Empire, and Arxiv-Year [7]. We summarize their statistics in Table 1. $\mathrm{homo\_ratio}$ represents the homophily ratio, a metric proposed by [42]. which is employed to gauge the degree of homophily within the graph. A lower $\mathrm{homo\_ratio}$ signifies a greater degree of heterophily, indicating a higher prevalence of edges that connect nodes of differing classes.
Table 1: Statistics of the datasets.
Dataset $N$ $|E|$ Feat. # Classes $\mathrm{homo\_ratio}$
Squirrel 5,201 217,073 2,089 5 0.22
Chameleon 2,277 36,101 2,325 5 0.23
Cora-ML 2,995 8,416 2,879 7 0.79
Citeseer 3,312 4,715 3,703 6 0.74
AM-Photo 7,650 238,162 745 8 0.83
Snap-Patents 2,923,922 13,975,791 269 5 0.22
Roman-Empire 22,662 44,363 300 18 0.05
Arxiv-Year 169,343 1,166,243 128 40 0.22
Table 2: Node classification results. Accuracy (%) with standard deviation for 10 runs. We highlight/underline the best/second-best method. For general GNN and non-local GNN baselines, we conduct experiments on both symmetrized versions and their directed counterparts, reporting better results from these two settings. OOM indicates out-of-memory.
Method Squirrel Chameleon Citeseer CoraML AM-Photo Snap-Patents Roman-Empire Arxiv-Year GCN 52.43$\pm$2.01 67.96$\pm$1.82 66.03$\pm$1.88 70.92$\pm$0.39 88.52$\pm$0.47 51.02$\pm$0.06 73.69$\pm$0.74 46.02$\pm$0.26 GAT 40.72$\pm$1.55 60.69$\pm$1.95 65.58$\pm$1.39 72.22$\pm$0.57 88.36$\pm$1.25 OOM 49.18$\pm$1.35 45.30$\pm$0.23 GraphSAGE 41.61$\pm$0.74 62.01$\pm$1.06 66.81$\pm$1.38 74.16$\pm$1.55 89.71$\pm$0.57 67.45$\pm$0.53 86.37$\pm$0.80 55.43$\pm$0.75 APPNP 51.91$\pm$0.56 45.37$\pm$1.62 66.90$\pm$1.82 70.31$\pm$0.67 87.43$\pm$0.98 51.23$\pm$0.54 72.96$\pm$0.38 50.31$\pm$0.42 MixHop 43.80$\pm$1.48 60.50$\pm$2.53 56.09$\pm$2.08 65.89$\pm$1.50 87.17$\pm$1.34 41.22$\pm$0.19 50.76$\pm$0.14 45.30$\pm$0.26 GPRGNN 50.56$\pm$1.51 66.31$\pm$2.05 61.74$\pm$1.87 73.31$\pm$1.37 90.23$\pm$0.34 40.19$\pm$0.03 64.85$\pm$0.27 45.07$\pm$0.21 GCNII 38.47$\pm$1.58 63.86$\pm$3.04 58.32$\pm$1.93 64.84$\pm$0.71 83.40$\pm$0.79 48.09$\pm$0.09 74.27$\pm$0.13 57.36$\pm$0.17 DGCN 37.16$\pm$1.72 50.77$\pm$3.31 66.37$\pm$1.93 75.02$\pm$0.50 87.74$\pm$1.02 OOM 51.92$\pm$0.43 OOM DiGCN 33.44$\pm$2.07 50.37$\pm$4.31 64.99$\pm$1.72 77.03$\pm$0.70 88.66$\pm$0.51 OOM 52.71$\pm$0.32 48.37$\pm$0.19 MagNet 39.01$\pm$1.93 58.22$\pm$2.87 65.04$\pm$0.47 76.32$\pm$0.10 86.80$\pm$0.65 OOM 88.07$\pm$0.27 60.29$\pm$0.27 DUPLEX 57.60$\pm$0.98 61.25$\pm$0.94 67.60$\pm$0.72 72.26$\pm$0.71 87.80$\pm$0.82 66.54$\pm$0.11 79.02$\pm$0.08 64.37$\pm$0.27 DiGCL 35.82$\pm$1.73 56.45$\pm$2.77 67.42$\pm$0.14 77.53$\pm$0.14 89.41$\pm$0.11 70.65$\pm$0.07 87.94$\pm$0.10 63.10$\pm$0.06 DirGNN 75.19$\pm$1.26 79.11$\pm$2.28 66.57$\pm$0.74 75.33$\pm$0.32 88.09$\pm$0.46 73.95$\pm$0.05 91.23$\pm$0.32 64.08$\pm$0.26 CGNN 77.61$\pm$1.87 79.54$\pm$1.82 70.27$\pm$0.66 77.06$\pm$0.67 90.41$\pm$0.32 72.33$\pm$0.16 92.89$\pm$0.51 66.97$\pm$0.15
6.3 Overall Results and Analysis
Table 2 reports the node classification results across eight digraph datasets. Our method CGNN achieves new state-of-the-art results on 6 out of 8 datasets, and comparable results on the remaining ones, validating the superiority of CGNN. We provide more observations as follows. Firstly, while non-local GNNs have the potential to cover the commute paths between adjacent nodes by stacking multiple layers, they do not consistently outperform general, shallow GNN models. It suggests that coarsely aggregating all nodes in commute paths is ineffective. The reason is that deeper models may aggregate excessive irrelevant information for the central node. Our goal is to encode mutual relationships between adjacent nodes by considering their commute times. Aggregating all nodes along the entire path introduces excessive information about other nodes unrelated to the direct relationship between the target nodes. Secondly, GNNs tailored for digraphs do not seem to bring substantial gains. Our results show that with careful hyper-parameter tuning, general GNNs can achieve results comparable to, or even better than, those tailored for digraphs, as evidenced in the Squirrel, Chameleon, and AM-Photo datasets. Thirdly, CGNN achieves state-of-the-art results on both homophilic and heterophilic digraph benchmarks. Notably, DirGNN also performs comparably on heterophilic graphs (e.g., Squirrel and Chameleon), supporting the findings of [7] that distinguishing edge directionality during message passing enables the central node to adaptively balance information flows from both heterophilic and homophilic neighbors, effectively mitigating the impact of heterophily. Moreover, CGNN, an enhanced version of DirGNN, further improves performance on these graphs by effectively incorporating commute times to refine the strength of relationships between nodes, enhancing model robustness under heterophily.
To illustrate this, we further examine the relations between commute-time-based proximity and label similarity along edges. As shown in Eq. 11, we use commute-time-based proximity $\widetilde{\mathcal{C}}$ to weigh the neighbors during neighbor aggregation. Then we define a label similarity matrix $\mathcal{M}$ where $\mathcal{M}_{ij}=1$ if $v_{j}\in\mathcal{N}_{i}$ and $y_{i}=y_{j}$; otherwise $\mathcal{M}_{ij}=0$. Essentially, $\mathcal{M}$ extracts the edges connecting nodes with the same classes from the adjacency matrix $\mathbf{A}$. Thus a higher value of $\|\mathcal{M}-(\mathbf{A}+\mathbf{A}^{\top})\|_{2}^{2}$ indicates a more pronounced negative impact of heterophily on the model’s performance. On the other hand, we compute $\|\mathcal{M}-(\widetilde{\mathcal{C}}^{\text{in}}+\widetilde{\mathcal{C}}^{\text{out}})\|_{2}^{2}$ to evaluate the efficacy of $\widetilde{\mathcal{C}}$ in filtering heterophilic information. The closer $(\widetilde{\mathcal{C}}^{\text{in}}+\widetilde{\mathcal{C}}^{\text{out}})$ is to $\mathcal{M}$, the more effectively it aids the model in discarding irrelevant heterophilic information. Table 3 visually demonstrates these relationships. We observe that in heterophilic datasets, the commute-time-based proximity matrix $(\widetilde{\mathcal{C}}^{\text{in}}+\widetilde{\mathcal{C}}^{\text{out}})$, aligns more closely with the label similarity matrix $\mathcal{M}$ than $(\mathbf{A}+\mathbf{A}^{\top})$. It indicates that $\widetilde{\mathcal{C}}$ effectively filters out irrelevant information during message passing by appropriately weighting neighbors, which explains the exceptional performance of CGNN on heterophilic datasets.
Table 3: The distance between $\mathcal{M}$ and $\mathbf{A}$, and between $\mathcal{M}$ and $\widetilde{\mathcal{C}}$.
(a) Squirrel
(b) AM-Photo
Figure 3: Accuracy vs. running time on Squirrel and AM-Photo.
Application scope analysis.
Can commute times always enhance message passing on directed graphs? To answer this, we analyze the scope of use for CGNN based on Table 2. For example, on the Snap-Patents and CoraML dataset, we observed that adding commute time-based weights during message passing did not significantly enhance performance. Now we can analyze the reason from the perspective of dataset. CoraML is a directed citation network where nodes predominantly link to other nodes within the same research area. However, in such networks, reciprocal citations between two papers are impossible due to their chronological sequence. Consequently, mutual path dependencies do not exist, and thus, incorporating commute times to adjust neighbor weights might might (slightly) hurt performance. A similar situation exists with the Snap-Patents dataset, where each directed edge represents a citation from one patent to another, again indicating the absence of mutual path dependencies.
In conclusion, in networks like citation networks where mutual relationships inherently do not exist, applying commute times to enforce these relationships is unnecessary. Conversely, our model is particularly effective in networks like webpage networks and social networks—examples being Squirrel and AM-Photo—where mutual relationships are prevalent. For instance, in a social network, an ordinary user may follow a celebrity, creating a short path to the celebrity. However, the reverse path from the celebrity back to the user might be considerably longer. Therefore, in such networks, considering mutual relationships based on commute times can provide a more accurate description of node relationships.
Table 4: Accuracy and running time (s) of CGNN and CGNNppr.
Squirrel Chameleon Citeseer CoraML AM-Photo Method Acc. Time Acc. Time Acc. Time Acc. Time Acc. Time CGNN 77.61 102.93 79.54 112.24 70.27 92.26 77.02 127.15 90.41 121.80 CGNNppr 68.37 257.84 71.69 253.05 68.59 137.82 76.23 192.09 88.52 203.04
6.4 Efficiency Comparsion
Fig. 3 compares the accuracy of CGNN and other baseline models with their running times. Despite the additional computational load of calculating commute-time-based proximity, the results show that CGNN provides the best trade-off between effectiveness and efficiency. In particular, on the Squirrel dataset, CGNN has the third-fastest calculation speed while yielding accuracy nearly double that of all other methods. On AM-Photo, CGNN achieves the highest accuracy while maintaining moderate efficiency.
6.5 Component Analysis
Table 5: Changes in commute times before and after rewiring.
CoraML Chameleon Roman-Empire $\delta$ 0.03280 0.00157 0.06242
Comparison between graph rewiring and PPR.
In Section 5.2, we construct a rewired graph $\widetilde{G}$ based on feature similarity to guarantee the irreducibility and aperiodicity. This approach introduces at most two additional edges per node, specifically targeting those with the highest feature similarity, while minimally altering the original graph structure to preserve semantic information. To investigate changes in commute times before and after rewiring, for the original graph, we use its largest connected component, removing absorbing nodes (i.e., nodes without outgoing edges) to ensure that we can compute meaningful and deterministic commute times. We denote the average normalized commute time for the original graph as $c_{orig}$; For the rewired graph, we directly compute the commute time and denote this average normalized value as $c_{rew}$. We then use $\delta=\frac{\|c_{orig}-c_{rew}\|_{2}}{\|c_{orig}\|}$ to quantify the changes, which can be intepretated as the proportion of commute information changed in the original graph. As shown in Table 5, the graph rewiring method can effectively preserve the original commute times of the graph.
In contrast, the classic PageRank transition matrix, defined as $\mathbf{P}_{pr}=\gamma\mathbf{P}+(1-\gamma)\frac{ee^{\top}}{N}$, achieves a similar objective but results in a completely connected graph $G_{pr}$. However, this approach tends to overlook the sparse structure of the original graph, which may alter the semantic information in the graph. Additionally, computing commute times using a dense transition matrix incurs a high computational cost. To validate the effectiveness of the rewiring approach over the PPR method, we conduct an experiment where $\widetilde{G}$ is replaced with $G_{pr}$ in the computation of commute-time-based proximity. We denote this variant as ‘CGNNppr’ and the results of accuracy and efficiency are reported in Table 4. The findings reveal that the PPR approach is suboptimal in terms of both accuracy and efficiency, thereby underscoring the effectiveness of our rewiring-based approach.
Table 6: Impact of directed structure.
Squirrel CoraML AM-Photo CGNN 77.20 77.02 90.41 CGNNsym 71.31 70.25 88.42
Directed vs. Undirected.
To validate the critical role of directed structures in our model, we transform all directed edges into undirected ones by adding their reverse counterparts. This process results in a symmetric adjacency matrix, denoted as $\mathbf{A}_{\mathrm{sym}}$. Subsequently, the commute time is calculated based on the transition matrix derived from $\mathbf{A}_{\mathrm{sym}}$. We refer this variant as ‘CGNNsym’. Table 6 shows the accuracy of CGNN and CGNNsym on three datasets. We find that edge direction can significantly influence the prediction accuracy for our model.
7 Conclusion
Identifying and encoding asymmetric mutual path dependencies in directed graphs is essential for accurately representing real-world relationships between entities. In this work, we utilize the concept of commute time to assess the strength of relationships in directed graphs and introduce the Commute Graph Neural Network (CGNN) to incorporate node-wise commute time into node representations. To achieve this, we propose DiLap, a novel Laplacian formulation derived from the divergence of the gradient of signals on directed graphs, along with an efficient computational method for deterministic commute times. By integrating commute times into GNN message passing through neighbor weighting, CGNN effectively harnesses path asymmetry in directed graphs, thereby enhancing node representation learning. Our extensive experiments across eight directed graph datasets demonstrate that CGNN significantly outperforms existing methods.